May 7, 2013 Calorimetry

From GlueXWiki
Jump to: navigation, search

'Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

  1. Light guides
    • Return 84 non-conforming light guides (Elton) Boxes are prepared and ready for shipment
    • Prepare letter for customs to prepare for exchange (Alison) Completed

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Monitoring: Media:BCAL_update_7_May_2013.pptx
  4. FCAL
  5. BCAL
  6. Reconstruction and Simulation
  7. TDR update
  8. Calibration of Calorimeters
  9. Any other business


Attendees: Elton, Beni, David, Yi (JLab), Matt (IU), Christina, George Vasileiadis (Athens), Alam, Sergey, Elias, Will (USM), Shaun, Andrei, Zisis (Regina)

  1. Announcements: None
  2. Action Items: Alison letter done. Elton is working on shipping LGs to USM.
  3. Monitoring: G. Vasileiadis presented new measurements on the FCAL Boards. The LED mini boards were binned into 5 groups of 10. Rationale was presented, having strong LEDs closest to the opposing inner corner (close to beam) in order to meet requirements of amount of light there. After discussion, it was decided that Christina would notify the CAL WG by email, giving folks a day to comment on the results, after which all 5 groups will be shipped to JLab. One (Athens recommends Group D) will be kept as a spare.
  4. FCAL: The installation is moving along. Referees comments on the NIM article were received with only minor issues raised connected to the energy resolution being consistent with previous operations. The FCAL contract was extended and will be sent soon.
  5. BCAL:
    1. Zisis has received no response from Jack McKisson yet on which SiPM can stay in Regina. It was suggested to to put a deadline (this Thursday) to the response.
    2. Shaun is looking at the dual-trigger counter data (one weekend long) on the BCAL module cosmics from the ESB tests. Very few events survive stringent cuts and Shaun will now loosen them. Elton reminded us of the urgency to get functional scripts and check the accumulated data from the various completed modules, to ensure that no issues exist.
    3. Elton reported LG gluing. No issues are apparent ever since the new procedure was implemented. These new modules will be tracked over time, so have not been crated. Three of the old procedure modules are still in the ESB and be uncrated and evaluated for repairs. Repairing/replacing LGs along the perimeter is relatively easy, and ideas were exchanged on how to fix interior ones. Removing their nearest neighbor along an edge may be the way to do this. Checking modules 12 hours after gluing (curing process) and 72 hours later (setting process) should verify whether they will be ok for the long run. During the upcoming Reconstruction Meeting (Thursday) we will discuss whether we can simulate effects of poor glue surfaces either with the full HDGEANT or a toy MC. It was noted that the airgap introduces a 10% loss of light (verified by Regina and Athens). Also, whereas sampling dominates the energy resolution, the z-coordinate location is dominated by photostatistics, so these cannot be disregarded.
    4. Elton describe problems with LGS, with a large fraction of them not meeting specs. These were classified in 3 groups: a) length issues, with some short as much as 0.7mm, b) leaners, which have surfaces that do not sit flat (perpendicularly) to the calorimeter faces, and c) oversized ones. Soon a decision has to be taken on how to handle these, and this may include shipping them vack to USM. 110 will be shipped back soon.
  6. Reconstruction and Simulation: Zisis reported that Will Levine checked in the IU code with separate treatment of ADC and TDC info, and first-pass time-walk corrections. Will, Andrei and Irina are woking on different aspects of this effort, and a report will be given at the Collaboration meeting.
  7. TDR update: Zisis finished and checked in the BCAL section. George Lolos is reviewing it, and once his comments have been incorporated Zisis will open it for comments from anyone.
  8. Calibration of Calorimeters: Elton reported on the JLab programmable pulsing system (VME 1495). Ben Raydo finished programming the FPGA and Hovannes implemented it in slow controls with nice GUIs. Tara Bogart will work on getting it going in the ESB area with long term goal of implementing it for the FCAL and BCAL.
  9. Any other business: None